one of the solutiin of the equation (1-x^2) y''-2xy'+12y=0 is
Ans-P3(x)
1
Expert's answer
2021-09-10T00:02:58-0400

Let "y=ax^3+bx^2+cx+d." Then


"y'=3ax^2+2bx+c"

"y''=6ax+2b"

Substitute


"(1-x^2) (6ax+2b)-2x(3ax^2+2bx+c)"

"+12(ax^3+bx^2+cx+d)=0"

"(-6a-6a+12a)x^3+(-2b-4b+12b)x^2"

"+(6a-2c+12c)x+2b+12d=0"

"x^3:0=0"


"x^2:6b=0"


"x^1: 6a+10c=0"


"x^0: 2b+12d=0"


"c=-\dfrac{3}{5}a, b=0, d=0"

"y=ax^3-\dfrac{3}{5}ax, a\in\R"



One of the solutiin of the given equation is


"y=5x^3-3x"